OpenStudy (anonymous):

multiple choice, State whether the following sequence is arithmetic; and if so, state the common difference. -1, -3, -5, -7, ... 1)2 2)-2 3)not arithmetic 4)-3

OpenStudy (anonymous):

Do the differences below all have the same answer ? (-7)-(-5)=? (-5)-(-3)=? (-3)-(-1)=?

OpenStudy (anonymous):

i have no idea!

OpenStudy (anonymous):

The 4th number minus the third number: (-7)-(-5) = -7 + 5 = 5 - 7 = -2 The 3rd number minus the second number: (-5)-(-3) = -5 + 3 = 3-5 = -2 etc. If you check all the numbers, you will see that the difference between each number in the sequence and the previous number in the sequence is always -2

OpenStudy (anonymous):

This means it is an arithmetic sequence, and the common difference is -2.

OpenStudy (anonymous):

oh!!!!! thankyou!